Michael owns 3/8 of the shares in a company . he sold 1/3 of his shares to his friend . if there are a total of 7600 shares in the company .how many share does Michael still have left ?
Found 2 solutions by jorel1380, MathTherapy:
Answer by jorel1380(3719)

If he sold 1/3rd of his shares, then he has 2/3rds left. So 2/3 x 3/8=2/8=1/4 of all shares of the company. 1/4 x 7600=?????
☺☺☺☺

Answer by MathTherapy(10552)


Michael owns 3/8 of the shares in a company . he sold 1/3 of his shares to his friend . if there are a total of 7600 shares in the company .how many share does Michael still have left ?
Remaining shares: